OpenStudy (anonymous):

Find the interval of convergence of:

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty}n4^{-n}(x-4)^{n}\]

OpenStudy (anonymous):

\[ \frac{a_{n+1}}{a_n}=\left | \frac{(n+1) 4^{-n-1} (x-a)^{n+1}}{n 4^{-n} (x-a)^n}\right | =\frac{(n+1) \left| x-a\right| }{4 n} \]

OpenStudy (anonymous):

\[ \lim_{n\to \infty} \frac{a_{n+1}}{a_n}=\lim_{n\to \infty}\left | \frac{(n+1) 4^{-n-1} (x-4)^{n+1}}{n 4^{-n} (x-4)^n}\right | =\lim_{n\to \infty}\frac{(n+1) \left| x-4\right| }{4 n}=\frac {|x-4|}{4}<1\\ |x-4| < 4 \\ -4< x -4< +4\\ 0< x <8 \] We have to examine the end points separately

OpenStudy (anonymous):

It is easy to see that the series does not converge for x=0 or x =4

OpenStudy (anonymous):

I'm still a bit confused, what is the interval of convergence then?

OpenStudy (anonymous):

I am confused where you had "does not converge for x=4", as arent we examining the end points 0 and 8?

OpenStudy (anonymous):

From my checking of endpoints, the series don't converge at the point x=0 and x=8. So the interval of convergence is 0<x<8 ?

OpenStudy (anonymous):

@eliassaab